Samer's weight on earth is 588 N. Using the law of universal gravitation, what will be his weight on Mars?


1
Expert's answer
2021-01-12T12:10:20-0500

"F=G\frac{Mm}{R^2}"


"F_e=G\frac{M_em}{R_e^2}" and "F_m=G\frac{M_mm}{R_m^2}" "\to"


"\frac{F_eR_e^2}{GM_e}=\frac{F_mR_m^2}{GM_m}\to F_m=\frac{F_eR_e^2M_m}{M_eR_m^2}="


"=\frac{588\cdot 6371000^2\cdot 6.39\cdot 10^{23}}{5.972\cdot 10^{24} \cdot 3376000^2}=224(N)" . Answer